Circular No. 7822

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V838 MONOCEROTIS T. Iijima, Padua Astronomical Observatory, Asiago; and M. Della

Valle, Arcetri Astrophysical Observatory, Florence, communicate: “Highdispersion spectra (range 430–690 nm, resolution 0.06 nm) of V838 Mon were taken on Feb. 1.92, 2.93, and 3.95 UT with the Padua/Asiago 1.82m telescope (+ échelle spectrograph). The spectral features drastically changed between Feb. 1 and 2. In addition to the emission lines of Na I and Ba II, which were visible in early spectra (IAUC 7785, 7786, 7812), many lines of ionized metals (Fe II, Ti II, Cr II, Mg II, Si II, etc.) emerged. Most of the metallic emission lines are flanked by P-Cyg-type absorption components, blueshifted by ∼ −120 km/s. Hα exhibits ‘narrow’ emission (FWHM ∼ 160 km/s) flanked by a P-Cyg absorption component blueshifted by −150 km/s. The ‘narrow’ component of Hα is superimposed on a ‘broad’ emission characterized by a FWZI of ∼ 2800 km/s. This peculiar object seems to be undergoing a normal nova explosion.”

Visual magnitude estimates: Feb. 2.983 UT, 7.8 (A. Kammerer, Ettlingen, Germany); 4.042, 7.3 (J. Garcia, Rama Caida, Argentina); 5.038, 7.1 (R. Y. Shida, Sao Paulo, Brazil); 5.740, 6.7 (L. Kiss, Piszkéstető, Hungary); 7.530, 6.8 (A. Pearce, Nedlands, W. Australia); 8.784, 7.3 (K. Hornoch, Lelekovice, Czech Republic).

SUPERNOVA 2002ap IN M74 S. Smartt, Institute of Astronomy, Cambridge; and P. Meikle, Impe-

rial College, London, write: “A spectrum (range 569–610 nm; resolution 0.08 nm) of SN 2002ap, obtained with the William Herschel Telescope at La Palma on Feb. 2.9 UT by C. Benn, shows absorption features of the interstellar-medium lines Na I (588.995 and 589.592 nm) shifted by +631 km/s. This matches the redshift of M74 and confirms that SN 2002ap is indeed associated with that host galaxy. We see no higher redshifted components in this spectrum (continuum signal-to-noise = 130), which suggests that SN 2002ap is not behind M74 in an unassociated galaxy.”

COMET C/2002 C1 (IKEYA-ZHANG) Visual m_{1} estimates: Feb. 3.74 UT, 8.0 (M. Meyer, Hohe Wurzel, Ger-

many, 20×100 binoculars); 5.13, 7.8 (C. S. Morris, Fillmore, CA, 20×80 binoculars); 7.53, 7.9 (A. Pearce, Nedlands, W. Australia, 20×80 binoculars); 8.42, 7.9 (D. A. J. Seargent, The Entrance, N.S.W., 25×100 binoculars); 9.40, 7.4 (Y. Nagai, Yamanashi, Japan, 20×100 binoculars).

2002 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green